### v2

# SoK: DARPA’s AI Cyber Challenge (AIxCC): Competition Design, Architectures, and Lessons Learned

Abstract
DARPA’s AI Cyber Challenge (AIxCC, 2023–2025) is the largest competition to date for building fully autonomous Cyber Reasoning Systems (CRSs) that leverage recent advances in AI—particularly large language models (LLMs)—to discover and remediate vulnerabilities in real-world open-source software. This paper presents the first systematic analysis of AIxCC. Drawing on design documents, source code, execution traces, and discussions with organizers and all finalist teams, we examine the competition’s structure and key design decisions, characterize the architectural approaches of finalist CRSs, and analyze competition results beyond the final scoreboard. Our analysis reveals the factors that truly drove CRS performance, identifies genuine technical advances achieved by teams, and exposes limitations that remain open for future research. We conclude with lessons for organizing future competitions and broader insights toward deploying autonomous CRSs in practice.
1 Introduction
Open-source software (OSS) underpins critical infrastructure, yet scaling vulnerability discovery and remediation remains challenging. DARPA’s AI Cyber Challenge (AIxCC, 2023–2025) addresses this by challenging teams to build fully autonomous Cyber Reasoning Systems (CRSs) that leverage large language models (LLMs) to discover and patch vulnerabilities in real-world C and Java projects. The final competition in August 2025 represents the largest-scale evaluation of autonomous vulnerability analysis to date: around 143 hours of fully autonomous operation, CRSs from seven finalist teams analyzed 53 challenge projects derived from critical infrastructure software, each equipped with $85K in cloud compute and $50K in LLM API credits.
Despite the competition’s completion, no systematic study has examined AIxCC’s design rationale, the technical approaches employed by participating teams, or the lessons that emerged from this large-scale competition. Such an analysis would benefit multiple communities: competition organizers designing future challenges, security researchers developing advanced vulnerability detection and patching techniques, and practitioners seeking AI-driven security solutions.
To fill this gap, we conducted a systematic study of the final competition (AFC), drawing on multiple primary sources: all seven finalist CRS codebases and whitepapers, the complete competition database (challenges, results, and execution traces) from organizers, and discussions with organizers and all finalist teams. Our analysis examines AIxCC from three perspectives: the design decisions that shaped the competition, the architectural and technical choices made by finalist teams, and competition results and their implications. Specifically, we address the following research questions:
RQ1: How is AIxCC designed to guide and evaluate AI-driven vulnerability discovery and patching?
RQ2: What architectural and technical approaches did finalist teams employ?
RQ3: What insights emerge from the competition results?
RQ4: What are the lessons and future directions?
To address these questions, our work makes the following contributions:
A systematic analysis of AIxCC’s competition design, covering task design, scoring, challenge construction, and iterative refinement across rounds.
A taxonomy of CRS architectures and techniques across all seven finalist teams, spanning PoV generation, patch generation, SARIF validation, and bundling.
In-depth result analysis revealing the gap between technique capability and real-world effectiveness, grounded in per-CPV evaluation against foundational baselines.
Lessons on translating competition outcomes to industry deployment and research value, along with directions for future competitions.
To facilitate future research, all data, experiments, and artifacts will be made publicly available upon acceptance, with some materials subject to DARPA’s release timeline (see Appendix B).
2 Background: AIxCC as Competition
Competition Brief. AIxCC [19] is a two-year competition (2023–2025) organized by DARPA and ARPA-H, in collaboration with major AI providers (Anthropic, Google, Microsoft, and OpenAI) and advisors from OpenSSF. The competition aims to advance fully autonomous vulnerability discovery and remediation for open-source software. From 42 teams, seven advanced via the semifinal (ASC, DEF CON 2024) [18] to the final (DEF CON 2025). All finalist CRSs, challenges, and infrastructure are being open-sourced [20].
Comparison with CGC. AIxCC is a spiritual successor to DARPA’s Cyber Grand Challenge (CGC, 2014–2016) [17], inheriting its core premise of fully autonomous CRSs. Key differences include: ➀ Scope: CGC focused on binary security, i.e., attack-defense of customized binaries on DECREE OS [66], such as crafting exploits and defensive patches against opponents. In contrast, AIxCC focuses on vulnerability discovery and remediation for real-world OSS development in C and Java. ➁ AI emphasis: AIxCC provides LLM infrastructure from AI providers, highlighting LLM-based technique exploration, a capability unavailable during CGC.
Final Competition (AFC). The final competition ran for around 143 hours, during which seven finalist teams deployed fully autonomous CRSs to analyze 53 challenge projects (CPs) in C and Java. CRSs operated without human intervention, managing their own infrastructure with $85,000 in Azure compute and $50,000 in LLM API credits per team.
Our Focus on AFC. This work focuses on the final competition (AFC), with brief discussion of the semifinal (ASC).111This artifact represents the authors’ own statements and does not constitute an official DARPA statement. Our data sources include: ➀ publicly released finalist CRS code and competition whitepapers; ➁ official competition database, complete challenge code, and vulnerability data obtained from the organizers, which are pending public release; ➂ organizer reflection documents; ➃ questionnaires and meetings with most finalist teams. For technical analysis (§ 6, § 7), at least two authors performed cross-validation. All artifacts will be open-sourced upon publication (some subject to DARPA’s release timeline).
3 Competition Design
3.1 Design Goal: Real-World Relevance
Figure 1 shows the final competition workflow. The competition is embedded in GitHub to simulate real OSS development, where development events trigger challenges for CRSs to analyze. Organizers designed four core CRS capabilities:
Full Scan: when maintainers tag a new release, detect and patch vulnerabilities across the full codebase.
Delta Scan: when new code is merged via pull requests, conduct targeted analysis on the incremental changes.
Report Synthesis: after analysis, correlate all above findings into per-vulnerability reports.
CRS Submissions. Organizers send multiple challenges (Full Scan or Delta Scan) during the competition, with some challenges including SARIF broadcasts for teams to review. The total score is the sum of all challenge scores, where each challenge score is computed from the following submissions:
PoV ( pts): an input triggering abnormal execution (e.g., crash) as proof of vulnerability.
Patch ( pts): a fix that resolves the vulnerability while preserving functionality.
SARIF assessment ( pt): a judgment on whether a SARIF report is valid.
Bundle ( pts): a grouping that links related findings for the same vulnerability.
Developer-Centric Scoring. Each submission type scores within a range due to time decay: immediate submissions earn full points while last-minute ones receive only half. The scoring weights reflect how much each submission benefits developers: ➀ Patch scores high as it directly resolves issues; ➁ PoV scores lower as it only demonstrates vulnerabilities; ➂ SARIF scores lowest as report validation is semi-subjective; ➃ Bundle rewards correct linkage but penalizes incorrect ones. Its actual score ranges based on how many finding types (PoV, Patch, SARIF) are linked and the score of each finding; a vulnerability with all three correctly linked can earn at most +7, while any incorrect linkage flips the reward into a penalty.
For inaccurate submissions, an accuracy multiplier penalizes the total challenge score. To encourage technique exploration while discouraging spam, the penalty is non-linear: low accuracy rates are penalized more severely (e.g., 90% accuracy negligible penalty; 50% 6% reduction; 40% 13% reduction). See scoring details in Appendix C.
Challenge Construction. All final competition challenges are based on OSS-Fuzz [4], providing build tooling and fuzzing harnesses as entry points for vulnerability detection. For project selection, organizers initially proposed a list of critical infrastructure and healthcare-critical OSS projects; then collaborated with the OSS-Fuzz team to onboard prioritized projects not yet supported. For vulnerability injection, most are manually crafted synthetics inspired by historical N-day issues, avoiding AI training data contamination. A small number are genuine 0-days discovered during development or sourced from unreported OSS-Fuzz findings.
3.2 Iterative Design
To handle the organizational complexity (multi-party collaboration, two-year timespan, AI evolution speed), organizers adopted an iterative approach rather than perfect pre-design.
Lessons from Semifinal. The ASC served as a litmus test, selecting 7 finalists from 42 competing teams. With fewer teams to manage, AFC relaxed infrastructure constraints: teams deployed their own CRSs on Azure rather than using a highly constrained sandbox. AFC also incorporated key lessons from ASC: ➀ ASC required patches to be submitted against self-discovered PoVs, disadvantaging CRSs that produced quality patches but struggled with vulnerability discovery; AFC addressed this via cross-team patch validation where patches are scored against all PoVs from all teams. ➁ Challenge adoption in ASC required significant engineering effort per repository; AFC streamlined this by adopting OSS-Fuzz.
Multi-Round Final Competition. AFC introduced a multi-round structure to progressively test both organizer infrastructure and team CRSs. Organizers first conducted multiple internal rounds from January to March 2025 to validate readiness and identify risks. Following these, three unscored exhibition rounds with increasing challenge counts and budgets were held before the scored final round. Each round mirrored the final environment, prohibiting teams from accessing their CRSs and requiring fully autonomous operation. These unscored rounds proved valuable for testing and debugging: both organizers and competitors identified issues and made adjustments that prevented failures in the final round.
4 Challenge Projects
LABEL:t:cp-overview presents the repositories, challenge projects (CPs), and challenge project vulnerabilities (CPVs) used in AFC. For AFC, organizers developed 48 CPs containing 63 CPVs, drawn from 24 OSS repositories. We exclude five CPs for which the organizer did not provide fuzzing harnesses, as they were unscored during the competition.
Repositories. The 24 repositories—14 C and 10 Java—span diverse application categories, from image-processing libraries to healthcare-critical software. The number of harnesses per repository varies considerably, from a single harness (dav1d, libavif, etc.) to 55 (ndpi), and repository sizes range from 16K (libexif) to 4.9M (wireshark) lines of code.
Challenge Projects. The 48 CPs comprise 16 full-mode and 32 delta-mode challenges, of which eight contain no injected CPV and serve solely as 0-day discovery targets. For delta challenges, CPV-containing delta commits are relatively compact, ranging from 163 (cc5) to 1,407 changed lines (av2), whereas no-CPV delta commits are substantially larger, from 3.2K (av3) to over one million lines (mg3). Project build times range from 16 s (mg3) to 492 s (ws1), and harness sizes from 2.4 KB (av3) to over 20 GB (ws1). See LABEL:t:cp-details for more details.
Challenge Project Vulnerabilities. The 63 CPVs—33 in full-mode and 30 in delta-mode—split into 40 C and 23 Java vulnerabilities, covering 34 unique CWE types including memory corruption, path traversal, and command injection. Organizers also issued 13 SARIF broadcasts —8 valid and 5 invalid— to test CRS triage capability. See LABEL:t:cpv-details and LABEL:t:sarif-broadcast-details in the appendix for details.
5 Cyber Reasoning Systems
Table I summarizes the background information behind each team’s CRS. While all teams built systems targeting the same CRS core capabilities (§ 3.1), their architectural approaches varied widely, shaped by team expertise, resource constraints, and strategic priorities. We briefly introduce each team’s design philosophy, providing context for understanding their technical choices in subsequent sections.
ID Team CRS Bg Lang‡ LLM Lib AT Team Atlanta Atlantis Mixed Py,Rust LangGraph§ TB Trail of Bits Buttercup Industry Py LangGraph TI Theori RoboDuck Industry Py,Rust Self-built§ FB Fuzzing Brain† FuzzingBrain Academic Py,Go § SP Shellphish Artiphishell Academic Py Self-built§ 42 42-b3yond-6ug BugBuster Academic Py,Go LangChain§ LC Lacrosse Lacrosse Industry Py,Lisp DSPy
Atlantis: Ensemble-First Design. AT is built around the ensemble philosophy [13, 25, 32]: any technique demonstrating unique contribution is worth incorporating, and combining multiple independent approaches enhances overall robustness. This leads to multiple independent bug-finding modules collaborating through seed sharing, and eight patching agents with diverse repair strategies.
Buttercup: Expertise-Driven Decomposition. TB leverages domain expertise to design deterministic workflows that decompose challenges into well-defined subtasks, with LLMs integrated only where traditional tools fall short. Notably, TB avoids high-end reasoning models, believing that well-decomposed problems paired with mid-tier models suffice.
RoboDuck: Agentic Design around Bug Candidates. TI embodies an agentic-first philosophy, building on a custom agent library that maximizes autonomous LLM operation. The entire system revolves around bug candidates: from identification and filtering, through PoV generation and patching, to SARIF validation and bundling.
FuzzingBrain: Simple Architecture, Diverse LLM Strategies. FB balances engineering effort against performance by simplifying architectural design while maximizing LLM strategy diversity. Notably, over 90% of its codebase is vibe-coded [31]. It implements 23 independent strategies, each as a standalone Python script with minimal dependencies, varying in scope, depth, and language-specific handling.
Artiphishell: Comprehensive Technical Coverage. SP achieves the most comprehensive technical coverage, implementing diverse techniques across all four core capabilities. To coordinate these techniques (53 components), the team built a custom orchestration platform that launches them on-demand and facilitates inter-component communication.
BugBuster: Pragmatic Technology Choices. 42 follows a pragmatic philosophy, preferring simple and stable technology choices. For bug finding, the design is traditional fuzzing and program analysis centric, with LLMs limited to auxiliary roles like seed generation. When adopting academic techniques [54, 72], the team consistently simplifies them to be practical, replacing sophisticated optimizations.
Lacrosse: DSPy-Based Multi-LLM Workflow. LC uses a Lisp-based task distributor to coordinate fuzzing [43, 42], patching, and analysis in a multi-agent system [44]. DSPy [58] manages diverse LLMs in parallel or as fallbacks, with patching failures feeding back to refine vulnerability characterization.
6 Taxonomy of CRS Techniques
6.1 PoV Generation
PoV generation is a typical vulnerability mining scenario: each project provides one or more fuzzing harnesses, and CRSs must find inputs that trigger bugs reproducibly through these harnesses. Depending on challenge type, vulnerabilities may exist anywhere in the repository (full-mode) or within a pull request’s diff (delta-mode).
Two Complementary Pipelines. Table II presents finalist teams’ exploration across all major technical directions (Table VIII in the appendix provides further details). Two complementary pipelines emerge: improving the fuzzing pipeline through traditional or LLM-enhanced techniques, and LLM-based PoV generation that directly leverages LLMs to identify vulnerabilities and generate exploit inputs. These pipelines reinforce each other: the fuzzing pipeline provides coverage information and valuable inputs to assist LLM-based generation, while LLM-generated results (even unsuccessful ones) serve as seeds boosting fuzzers’ effectiveness.
AT TB TI FB SP 42 LC Pre-Comp Corpus ● ○ ● ● ○ Seed Gen Agent ● ● ● ● ● ● Bootstrap ● ● ● ● ● Solve cov blocker ● ● ● ● Mutator/generator ● Grammar-aware ● ● ● Semantic Feedback ● Improved Sanitizer ○ ○ Dict Gen ● ○ ○ ○ Concolic Fuzzing ○ Directed Fuzzing ○ ○ Parallel Fuzzing ○ ○ ○ ○ ○ ○ ○ Corpus sync ○ ○ ○ ○ ○ ○ Added C fuzzers ○ ○ ○ ○ Fuzzing Added JVM fuzzers ○ Bug Cand. I.D. ● ● ● ● ● Cand. filter ● ● ● ○ ● Non PoV Gen usage ✓ ✓ ✓ ✓ PoV Gen Agent ● ● ● ● ● W/ CWE guidance ● ● ● ● LLM Based PoV Gen Reach-then-exploit ● ● ● LLM PoV Gen Fuzz ✓ ✓ ✓ ✓ ✓ Pipeline Co-op Fuzz LLM PoV Gen ✓ ✓ ✓ Deduplication ○ ○ ● ● ○ ○ ○ PoV Sub. ASAP Submission ✓ ✓ ✓ ✓ ✓ ✓ ✓
Improving Fuzzing Pipeline. Most teams (5/7) simultaneously explored both pipelines, while 42 and LC focused exclusively on fuzzing. Among the five, AT and SP conducted the most comprehensive exploration; TB and TI selectively explored techniques directly targeting fuzzing seeds (reuse, generation, sharing), and FB minimized fuzzing exploration (only parallel fuzzing).
Pre-competition corpus. Five teams reused pre-collected corpora to bootstrap fuzzer initial coverage, typically with two steps: ➀ collecting and grouping seeds from public databases (ClusterFuzz [29], OSS-Fuzz [4], GitHub, etc.) before competition; ➁ matching relevant seeds to each harness via coverage-based ranking or similarity-based selection using harness names or LLM-inferred formats.
LLM-based seed generation. Teams use LLMs to generate seeds in two scenarios: early-stage bootstrap (analyzing harness code for input formats) and troubleshooting (generating inputs for coverage blockers [62]). Teams typically built agents for these scenarios, incorporating conventional program analysis tools for better performance. Interestingly, despite the goal being harness input generation, all teams chose to have LLMs generate Python scripts that produce inputs upon execution. Additionally, AT, TI, and SP explored generating input generators/mutators and explicit grammars to further enhance input quality.
LLM-Based PoV Generation Pipeline. Six teams explored this pipeline. The typical workflow involves two steps: ➀ identifying and filtering bug candidates; ➁ generating PoVs targeting these candidates. LC performs bug candidate identification but uses results for no-PoV patch generation rather than PoV generation. TB skips explicit bug identification, allowing the agent to autonomously judge during PoV generation.
Bug candidate identification. Teams design agent systems that combine LLMs, static analysis tools, and predefined lists of likely vulnerable functions (sinks) to identify and filter bug candidates. Besides standard agentic component construction, two techniques are explored: TI provides logprobs-based classification [46] as an agent tool, enabling agentic classification with token-efficient confidence estimation; SP and LC employ multi-source voting [67], aggregating weighted scores from multiple tools and LLMs to rank high-potential candidates.
PoV generation. Similar to seed generation, teams construct agents leveraging static/dynamic analysis tools, but targeting crash-triggering inputs rather than coverage extension. Teams use CWE information [41] to guide agents; some decompose PoV generation into reachability and exploitability, building separate reach and exploit agents.
Pipeline Cooperation. Most teams explore cooperation between pipelines in both directions. On one side (LLM PoV Gen Fuzz), successful, failed, and intermediate results from LLM PoV generation are shared with fuzzers, expecting fuzzers to extend coverage or mutate these near-solutions into actual PoVs. On the other side (Fuzz LLM PoV Gen), fuzzers provide coverage information to guide LLM generation; for teams with separate exploit agents, reached-but-not-exploited inputs discovered by fuzzers are also shared for exploitation attempts.
Challenge-type-specific Adaptations. Overall, teams adopt a unified technical design, and most techniques explored are theoretically general. However, adaptations for different languages and challenge types can be necessary for better effectiveness.
Language-specific optimizations address vulnerability pattern differences: Java’s logical vulnerabilities often stem from unsafe sink function usage, prompting teams to adopt sink-targeted PoV generation, directed fuzzing toward sinks, and enhanced sanitizers.
For delta-mode challenges, teams narrow scope from full codebase to the diff and related code. For SARIF broadcasts received mid-competition, teams treat them as pre-specified bug candidates and incorporate these hints into the pipeline.
PoV Submission. All teams adopt straightforward strategies: submit unique PoVs as soon as possible. This simplicity stems from the scoring rules: correct but duplicate submissions incur only time-decay penalties, not accuracy penalties, making early submission always preferable. To minimize redundant submissions, all teams implement deduplication using crash stack traces, input hashing, sanitizer signatures, etc. TI and FB further use LLMs to group semantically equivalent PoVs.
6.2 Patch Generation
All CRSs follow a de facto patch pipeline as shown below, where RCA denotes root cause analysis and brackets indicate optional steps. Within this pipeline, teams explore different agent designs to build effective LLM-centric patch systems. Table III analyzes teams’ design choices across agent architecture and each pipeline step.
loop([RCA] → Generate → Validate) → Dedup → Submit
AT TB TI FB SP 42 LC Multi-Arch ● ● Multi-Agent ● ● Agent Design Single-Agent ● ● ● Standalone RCA ● ● ● ● Multi-PoV RCA ● ● ● Non-LLM RCA ● Contextualization† ● ● ● ● ● ● ● Code Indexer ● ● ● ● ● SAST ● ● ● ● CWE Guidance ● ● Fine-tuned LLM ● Agentic Code Search ● ● ● ● ● ● ● Dynamic Info ● ● ● ● PoV Bytes ● ● ● LLM Reflection ● ● ● ● ● Patch Gen No-PoV Patch ● ● ● Validity Checks ● ● ● ● ● ● ● Build ● ● ● ● ● ● ● PoV Test (Gen) 1 N 1 N 1 PoV Test (Submit) N N 1 Proj. Tests ● ● ● ● ● ● LLM as Judge ● ● ● Post-patch Fuzz ● ● Patch Valid. Rebuild Optim. ● ● Min. Patch Set Calc. ● ● ● ● Patch Sub. No-PoV Delayed Sub. – – ● ● – – ●
Patching Agent Design. CRSs’ agent designs are three-fold:
Multi-Arch: Ensembling multiple architectures improves both performance and reliability. AT ensembles diverse patcher agent architectures [32], believing that different agent designs excel in different cases. The eight agents span diverse designs: workflow-based pipelines with ReAct-style [71] tool use, autonomous agents with iterative context retrieval, multi-agent systems for handling context limitations, and off-the-shelf coding agents (Aider [26], SWE-Agent [70]). SP takes a different ensembling approach, combining a fully agentic LLM patcher with a program-analysis-assisted, one-shot minimal LLM patcher. When generating, AT stops when any agent successfully produces a valid patch, while SP collects all candidates and ranks them by submission strategy.
Multi-Agent: Multi-agent designs enhance LLM capability through agent specialization. TB decomposes patching into sub-problems with specialized agents for RCA, patch writing, context retrieval, and reflection. TI implements a ReAct-style nested agent loop, with SourceQuestionsAgent as a tool for code understanding, separating code exploration from patch generation.
Single-Agent: Three CRSs adopt single-agent designs with varying degrees of agentic customization. FB implements 23 strategies spanning delta-scan and full-scan modes, differentiated by context scope and knowledge injection. 42 maximizes configuration diversity, exploring 16 combinations of temperature and prompt context (failed cases, stack traces, etc) within a single agent architecture. LC uses a DSPy-based workflow with model escalation from cheaper to expensive models.
Patch Generation. Beyond agent architecture, CRSs employ various techniques to improve patch generation quality:
Standalone RCA: Four CRSs (AT, TB, TI, SP) implement standalone RCA components, allowing LLMs to separately focus on root cause analysis and patch synthesis as distinct sub-problems. All four build LLM agents for RCA, while SP additionally incorporates a non-LLM RCA component that combines multi-source signals (SAST reports, stack traces, fuzzing invariants, etc.) with weighted voting to rank root cause candidates.
Contextualization: CRSs augment patch generation with diverse information sources, including static analysis reports, dynamic execution traces, PoV bytes, project code, and CWE-specific guidance. Teams vary in their combinations, while AT has the most technique coverage due to its patch agent ensembling strategy. Further, AT is the only team that explored fine-tuned LLMs for context retrieval.
LLM Reflection: LLM reflection [55] enables agents to learn from failed attempts, and most CRSs have adopted it. One typical example is TB, which implements a dedicated reflection agent that analyzes failures at each generation step and provides corrective guidance.
No-PoV Patch: Three CRSs (TI, FB, LC) attempt patch generation without PoV. Without dynamic evidence, this approach carries risk, but can address vulnerabilities that are obvious to identify yet difficult to trigger with a PoV. The generation technique is similar to but more limited than PoV-based techniques, while all teams focus on risk mitigation by limiting no-PoV patch quantities per challenge and imposing stricter submission conditions (e.g., delayed submission, gated by prior success rate, etc).
Patch Validation. As shown in Table III, CRSs employ various validation checks to ensure patch correctness:
Basic Checks: PoV tests, build verification, and project test suites form the basic validation strategies across all CRSs (42 skips project tests).
LLM-as-Judge: Three CRSs incorporate LLM-based evaluation [30], including judging whether patches correctly address the root cause and follow the prescribed fix strategy (AT), and self-reflecting on whether patches genuinely fix vulnerabilities rather than being superficial, easily-bypassed, or having side effects (FB/SP for No-PoV/all patches).
Post-patch Fuzzing: FB and SP adopt short-term fuzzing on patched projects for incomplete patch detection.
PoV Optimization: Many CRSs use partial PoVs instead of all available PoVs to accelerate patch generation iteration. AT and FB use single PoV during generation then validate against more before submission, while TB/LC use multiple/single PoVs throughout, respectively.
Patch Submission. The simplest strategy is to submit a patch immediately for each deduplicated PoV, but this risks duplicate submissions and imperfect patch penalties. Teams adopt the following optimizations:
Minimal Patch Set Calculation: Four CRSs (AT, TB, SP, and 42) leverage the fact that a single patch can fix multiple PoVs sharing the same root cause, and compute a minimal patch set that covers all known PoVs to avoid duplicate submissions. These CRSs differ in three dimensions: ➀ Calc. timing: on each new PoV (AT, TB, SP) for faster response, or hourly (42) for better global optimization. ➁ Calc. mode: incremental over uncovered PoVs only (AT, 42) for simplicity, or recompute over all PoVs (TB, SP) for better optimization at the risk of duplicate submissions. ➂ Submission timing: immediate (AT, TB, 42), or delayed (SP, 60min) for better global minima.
No-PoV Patch Delayed Submission: All three CRSs with No-PoV capability (TI, FB, LC) delay submission to reduce imperfect patch penalties: TI waits 45min and gates on PoV patch success history, FB waits until 50% of challenge time, and LC submits 30min before deadline.
Challenge-type-specific Adaptation. With agentic architectures, CRSs primarily adapt to different challenge types by adjusting language-specific tooling and prompts, as well as CWE or bug-type-specific guidance. For delta mode, most CRSs focus vulnerability search on modified code from diff files. Additionally, TI runs two diff analysis agents in parallel, one filtering out compiler-unused files and one using the complete diff, to broaden vulnerability search from immediate change sites to all affected code.
6.3 SARIF Validation
CRS Category Submission Strategy Overview AT, TB PoV-centric FB PoV-centric TI Bug-cand-centric SP, LC LLM-judge-centric 42 LLM-judge-centric
The SARIF validation task requires CRSs to assess each static analysis report as valid or invalid, submitting a verdict of Correct or Incorrect. CRSs can resubmit to revise verdicts while incurring penalties in both time and accuracy.
Key Evidence for Validation. Table IV summarizes each team’s core submission strategy. Teams primarily relied on three types of evidence: ➀ Match Any PoV: matching SARIF locations against crash information from exploited vulnerabilities; ➁ Match Bug Cand.: matching any bug candidate (inferred from LLM, static analysis, PoV, etc.); ➂ LLM As Judge: agentic prompting to directly assess its correctness.
Validation Strategies. Teams adopted three strategies:
PoV-centric: AT, TB, and FB primarily rely on PoV matching, submitting Correct only when a match is found and withholding unmatched reports. FB additionally uses a fallback LLM judgement, but only submits Correct from it.
LLM-judge-centric: SP, 42, and LC rely on LLM judgement, submitting both Correct and Incorrect based on the model’s assessment. 42 falls back to PoV matching when the model replies with uncertainty.
Bug-cand-centric: TI matches SARIF reports against its bug candidate database, initially submitting Incorrect for unmatched reports and revising to Correct when new evidence emerges.
6.4 Bundling Strategy
Bundling pairs PoVs, patches, and SARIF assessments into coherent vulnerability reports. Unlike other submissions with time decay, bundling allows free updates until the deadline, with scoring based solely on final results. A bundle can contain any two of three pairings, PoV-Patch, PoV-SARIF, and Patch-SARIF, to form a complete scoring bundle, while any incorrect pairing will penalize the entire bundle.
Pairing Source AT TB TI FB SP 42 LC PoV-based Patch ● ● ● ● ● ● ● PoV-Patch Match No-PoV Patch ● SARIF Validation ● ● ● ● PoV-SARIF SARIF-guided PoV ● ● Bug Candidate DB ● Patch-SARIF SARIF-guided Patch ●
Table V summarizes each team’s bundling pairing strategies. Given the risk of score penalties, teams tend to derive pairings from existing workflows rather than inferring relationships independently. ➀ All teams naturally derive PoV-Patch relationships from PoV-based patch generation. For No-PoV patches (§ 6.2), TI retroactively links PoVs once discovered, while FB does not. ➁ For SARIF pairings, teams either reuse their SARIF validation results (§ 6.3) or use SARIF reports to generate PoVs/patches, pairing them upon success. 42 and LC do not participate in SARIF pairing, while only teams with No-PoV patch capability can submit Patch-SARIF bundles.
7 Competition Result Analysis
AT TB TI FB SP 42 LC C 52.6 31.0 22.7 22.6 31.4 49.1 1.5 Java 27.0 21.3 31.6 29.7 16.5 21.1 0.0 PoV Sum 79.6 52.4 54.3 52.3 47.8 70.1 1.5 C 113.5 74.2 51.8 45.3 40.5 9.7 4.9 Java 57.5 26.7 49.8 23.5 13.8 4.5 0.0 Patch Sum 171.0 100.9 101.6 68.8 54.3 14.2 4.9 C 5.0 1.0 3.9 4.7 7.5 8.7 0.0 Java 1.0 0.0 1.0 1.5 1.0 1.0 0.0 SARIF Sum 6.0 1.0 4.9 6.2 8.5 9.7 0.0 C 99.6 49.4 26.0 27.6 18.3 7.8 3.2 Java 36.6 15.7 23.8 -1.1 7.0 3.2 0.0 Bundle Sum 136.2 65.1 49.8 26.4 25.3 11.0 3.2 C 270.6 155.6 104.4 100.2 97.6 75.4 9.6 Java 122.1 63.8 106.3 53.5 38.3 29.7 0.0 Pen. -0.4 -0.6 -16.3 -13.5 -0.1 -0.3 -1.1 Total Final 392.8 219.4 210.7 153.7 135.9 105.0 9.6
7.1 What Scores Reveal (and Conceal)
Figure 2 presents the score accumulation over time, and Table VI provides a detailed breakdown. The competition spanned 142.7 hours across seven phases (P1–P7), with tasks released concurrently within each phase.
Stability Wins. AT emerged as the winner with 392.8 points, nearly 80% more than second-place TB (219.4). This dominance was primarily accumulated through sustained availability across all phases. Although AT already had a visible lead by late P2, the majority of its advantage came from subsequent phases where close competitors had stopped functioning. Both TB and TI were competitive early on, but plateaued after P3 and P4 respectively (telemetry shows no further activity after these points). LC successfully generated over 1,200 PoVs for one CPV before crashing due to resource exhaustion. Beyond those system-wide failures, we could also identify certain individual challenge-level failures among CRSs due to bugs: AT was unavailable on all poi Java challenges, TI may have had issues with the wireshark C challenges, 42’s low patch score (14.2) resulted from a bug in their patch submission system [61], etc.
Accuracy Matters. The gap between second and third is as close as 8.7 points (see zoomed inset in Figure 2). TI’s pre-penalty score was actually higher than TB’s, but the point accuracy penalty, the largest among all teams, dropped them to third.
Comparison on Score Breakdown. Though AT led in most individual scoring categories, the observations above highlight a fundamental challenge in interpreting score breakdowns: stability issues confound capability comparisons, as missing scores may reflect system failures rather than technical performance differences. Java scores are more meaningful: all Java CPVs were released in P1 and P2 (except log4j in P6), when most systems were still active. Among Java results, TI achieved the highest PoV score (31.6), demonstrating the strongest vulnerability discovery capability; AT led in patching (57.5) and bundling (36.6). 42 achieved the highest SARIF score (9.7); and SP showed the lowest accuracy penalty ().
7.2 Auxiliary CPV Annotation
Challenges in Per-CPV Analysis. Figure 3 shows teams’ per-CPV scoring data. However, understanding why a CRS succeeded or failed on a specific CPV is usually not feasible. Competition telemetry—traces, logs, and events submitted to the infrastructure—is inherently incomplete and unordered: teams selectively uploaded logs for performance reasons, and were not required to guarantee logical ordering under massive parallelism across their cloud clusters. Furthermore, even with complete logs, CRS architectures ensemble multiple techniques such as fuzzers, symbolic execution, LLM-based analysis, and heuristics, making attribution to any single technique inherently difficult.
CPV Annotation via Foundational Techniques. To enable deeper analysis of per-CPV performance, we evaluate foundational vulnerability discovery and patch generation techniques on AFC challenges as analytical aids. The rationale is that their performance under ideal laboratory conditions reflects the inherent difficulty characteristics of CPVs. This allows us to contrast which CPVs the finalist CRSs have solved against the performance of these foundational techniques. Note that unlike building end-to-end CRSs as baselines, we evaluate specific techniques on isolated tasks under ideal conditions for CPV annotation. These techniques do not need to handle the challenges of fully autonomous systems, such as resource allocation across up to 50 harnesses per CP, patch generation from self-discovered PoVs, deduplication, scheduling, etc.
PoV Annotation (PF). Parallel fuzzing is the de facto standard for automated vulnerability discovery. We use atlantis-multilang-given_fuzzer from OSS-CRS [57], evaluating only vulnerable harnesses with 16 cores per harness, shared-memory seed sharing, and OSS-Fuzz configurations used in AFC. Fuzzers run up to six hours or until all expected CPVs are found. To mitigate randomness, a CPV is marked fuzz-discoverable (gray cell in Figure 3) if any of the three runs succeeds. The evaluation takes 8,906 CPU hours.
Patch Annotation (MR and CC). MultiRetrieval (MR) [60] is a minimal security-specialized patch agent from OSS-CRS [57], with a basic loop structure (see § 6.2), no root cause analysis, relying on AST-based code search [64] and patch-specialized prompts. Claude Code (CC) [2] is a general-purpose coding agent without domain specialization. Both receive ground-truth PoVs with sanitizer logs directly and use claude-3-7-sonnet-20250219 (CC version 1.0.88, 2025-08-21). CC tries at most 3 patch generations per CPV; MR determines iterations autonomously. Each agent runs three times per CPV, and results are reported as the union across runs. Generated patches are first validated automatically (build, PoV reproduction, and functional tests), then manually verified for correctness by two independent reviewers, as patches passing automated checks can still be semantically incorrect [33]. In Figure 3, ● indicates a valid patch is generated; indicates at least one patch passed automatic checks but failed manual verification; and blank indicates no patch passed automatic checks. Total LLM cost is $119.53 (CC) and $221.69 (MR), and manual validation takes two person-weeks.
CRS Performance by Annotations. First, the annotation results reflect that AFC challenges do not primarily reward solving exceptionally difficult problems: around half are solvable by a single foundational technique (PF: 34/63 PoV; MR: 31/63 patches; CC: 33/63 patches). Combining these annotations with CRS performance, two contrasting patterns emerge. Most CRSs demonstrate genuine improvements beyond foundational techniques, finding and patching many CPVs that PF/MR/CC cannot: particularly in P1–P2 C and Java challenges where LLM-based reasoning proves essential. Yet, CRSs also unexpectedly underperform on annotated solvable challenges. The causes are multifaceted: system-wide failures (TB, TI ceased after P3–P4), critical bugs (42’s patch submission issue, SP’s configuration issue, AT’s Java poi failure, etc), and the long tail of real-world edge cases that automated systems cannot generically handle. This leads to an interesting observation: a CRS that reliably applies foundational techniques in real-world conditions would rank among the top three. Besides, looking back at Figure 3, AT’s dominance was largely accumulated from P3 to P7, precisely where other teams unexpectedly underperformed relative to what the annotations suggest should be achievable.
7.3 PoV Generation Analysis
PF Solvable CPVs. PF marks 34 of 63 vulnerabilities (54%) as fuzzer-discoverable, with notable language disparity: 30/40 C (75%) versus only 4/23 Java (17%). This gap stems from multiple factors: inputs for Java CPs involve richer semantic constraints (e.g., XML), AFC Java CPVs include many timeouts and OOMs that are inherently fuzzer-unfriendly, and OSS-Fuzz default seeds for Java targets are lower quality than those for C. Interestingly, for C challenges, wireshark (ws*), shadowsocks (ss*), and sqlite (sd*) contain mostly fuzzer-discoverable CPVs by design: the organizer’s challenge documents note these were intended to evaluate patching and deduplication rather than PoV generation difficulty. These challenges were released in P3–P7, implying a stability requirement for CRSs to accumulate points from them.
CRS over Parallel Fuzzing. CRSs solved 8 C and 14 Java vulnerabilities that PF could not, demonstrating substantial improvement over parallel fuzzing alone. This advantage can be largely attributed to LLM components. According to § 6.1, PF represents parallel fuzzing with corpus synchronization, which nearly all CRSs incorporate as a baseline capability. While non-LLM fuzzing enhancements were explored by only a few teams, improvement over PF is observed across most teams, suggesting LLM-based techniques as the primary contributor. FB provides clear evidence: beyond parallel fuzzing, they rely solely on LLM-based generation, yet they solved multiple CPVs that PF missed. CRSs uniquely demonstrate the following capabilities over PF:
Strong targeted detection capability. Delta challenges dominate non-PF-solvable CPVs solved by CRSs (15/22). Given diff-based hints about vulnerability locations, CRSs can analyze code changes and generate triggering inputs directly. Representative cases include fp202, mg101 (C), and cc103, po105 (Java). Notably, several such CPVs involve indirect calls (cu206, cu403) where function pointers obscure the path to vulnerable code, indicating that CRSs have advanced reasoning capabilities to overcome these indirections.
Overcoming input grammar obstacles. Some CPVs require inputs conforming to complex grammars that random mutation cannot satisfy. Examples include curl protocol exploitation requiring TLV-formatted inputs [15] (cu206), and structured file formats like PDF with embedded XFA (pb101) or XLSX for SSRF (po105). CRSs can comprehend format specifications from code and construct syntactically valid inputs.
Solving logical constraints. Some CPVs guard vulnerable paths with constraints that defeat fuzzer feedback mechanisms: regex patterns (cc103, po106), encoding transformations like URL encoding (cc406), Unicode normalization (cc607, po102), or zlib compression (pb102), symlink-based path obfuscation (cc708), and mathematical guards (cc510, tk610). LLMs can reason about these transformations and generate constraint-satisfying inputs.
PF-Solvable CPVs Missed by CRSs. Some CPVs solvable by PF were missed by CRSs. Our analysis identified infrastructure implementation limitations as the primary cause: trivial yet critical obstacles in handling real-world complexities, rather than fundamental detection capability gaps:
Broken system dependencies: Some CPs deviate from standard OSS-Fuzz setups, breaking CRS initialization. av202 uses FuzzTest with its own fuzzer instantiation; pdfbox CPs download their own jazzer, bypassing CRS’s pre-installed tools.
Heavy build process: Unexpected resource demands can exhaust CRS nodes. Typically, wireshark (4.9M LoC) produces 47 harnesses of 455 MB each, causing disk exhaustion or OOM.
Reproduction behavior mismatch: CRSs’ reproduction criteria can differ subtly from organizer criteria. Some stateful bugs require multiple executions to trigger: sd105 crashes on the second execution, and the organizer runs 100 times by default. Thus, CRSs verifying with a single execution discarded valid PoVs.
Crash deduplication granularity: ss100–ss104 are distinct heap-buffer-overflows in json_parse_ex, differing only by line number. Coarse-grained deduplication (libfuzzer tokens, function-level grouping, similarity) failed to distinguish them as separate vulnerabilities.
CPVs Unsolved by Both PF and CRSs. A subset of CPVs remains unsolved by both PF and all CRSs, offering insights for methodological improvements. Some require complex reasoning beyond current CRS capabilities: cu307 and cu408 involve cryptographic transformations (XOR/shift computations, AES encryption with Base64 encoding) that are prone to fail due to continuous LLM hallucination and require tool-assisted verification. A particularly challenging case emerges from the pdfbox ExtractTextFuzzer harness, which embeds four timeout and two OOM CPVs (pb103 to pb108). This configuration exposes two intertwined challenges: ➀ timeout bugs produce hard to distinguish crash signatures, making fine-grained deduplication difficult; ➁ shallow bugs block discovery of deeper ones: once pb107 triggers, repeated timeouts prevent exploration of pb105 and pb106. Given the complexity and risk involved, no CRS attempted on-the-fly patching or fine-grained timeout deduplication, resulting in minimal CPV coverage on this harness.
7.4 Patch Generation Analysis
MR/CC Patchable CPVs. MR and CC annotate 31 and 33 of 63 CPVs as patchable, respectively, with 36 unique CPVs covered in total. Although more than half of all CPVs are covered, this does not indicate superior patching capability; rather, it reflects that MR and CC can solve basic fix challenges where the root cause is directly surfaced by the sanitizer report and the local code context suffices to derive a fix without cross-function or cross-file reasoning: textbook vulnerabilities with well-known fix patterns. Typical examples include boundary checks for buffer overflows (fp202, ss100–ss104), secure XML parser configuration for XML External Entity (XXE) (pb100, pb101), decode-then-normalize reordering for path traversal (cc406, cc607), single-line fixes such as null checks and format-string corrections (cm101, cu501), and CTF-style backdoors labelled by a “flag” string that simply need deletion (cu206, sd103, cc510).
Semantically Incorrect Patches from MR/CC. A notable fact is that a significant fraction of generated patches pass all automatic validation, yet contain semantic issues caught only by manual review (CC: 20/53, 37.7%, MR: 26/57, 45.6%).
Wrong root cause. Agents may suppress the crash symptom rather than addressing the underlying defect. One typical pattern is defensive patching in parsing logic, when root cause and crash location are distant (av202, sd105, pb103, pb104). For timeout CPVs (pb103, pb104, po106, pb108), a common error is inserting a hard-coded iteration limit instead of fixing the infinite loop root cause.
Incomplete fix. Agents can fix the specific crashing path instead of completely remediating the underlying bug. In mongoose (mg100, mg101, mg202), #line directives hide the amalgamated compilation unit from the sanitizer reports, so agents patched the reported files but missed it. Other cases include hardening an XXE parser but omitting a critical security setting (pb100), nullifying a dangling pointer at the call site instead of inside the buggy memory API (sd105, ex201), etc.
Functionality deviation. Some patches eliminate the crash but subtly alter program semantics in ways that functional tests do not cover (e.g., ws204, ws102, ws105, ws111, lx304, sd101, tk610, fp303). In av202, a patch produces incorrect edge colors because YUV-to-RGB conversion depends on neighboring pixels that the patch mishandles. In cc708, a patch resolves relative symbolic links to absolute paths during archive extraction, when it should only validate against path traversal without modifying the link target.
Introducing new bugs. In ss100, an extra break renders part of the encoding logic unreachable; in sd105, removing a mfree call introduces a memory leak.
Missing domain knowledge. In ws713, correct patching requires understanding how GVCP [1] bootstrap registers, which are standardized memory-mapped addresses for device discovery, are managed within the project. In ws101, patches confused self-reported packet lengths with verified ones, lacking the knowledge to recognize which length value can be trusted.
CRS over MR/CC. On one hand, CRSs solved 16 cases where MR/CC failed, demonstrating capabilities beyond foundational agents, such as the correct fixes for incomplete remediations (sd105, ex201), functionality deviations (ws111, lx304, sd101, tk610), and misleading sanitizer reports (mg202, mg100) listed in the previous paragraph. On the other hand, CRSs also suffer from semantically incorrect patches, though mostly at lower rates than MR/CC (see Table XII in Appendix M). AT and TB achieve 83.8% and 79.2% patch accuracy, respectively; SP reaches 100%, though we attribute this to a combination of strict patch filtering (e.g., its 5-minute post-patch fuzzing step) and patched challenge distribution, rather than concluding it fully mitigates semantic correctness issues. In contrast, TI and FB fall to 31.7% and 23.3% accuracy, concentrated in bursts of invalid submissions on one or two challenges, likely tied to implementation-specific strategies such as parallel generation without adequate deduplication. Despite these differences, a common factor among higher-accuracy CRSs is the adoption of multi-PoV validation, post-patch fuzzing, or LLM-based reflection (§ 6.2), yet incorrect patches remain prevalent overall.
MR/CC Patchable CPVs Failed by CRSs. Nine CPVs that MR or CC patch under laboratory conditions were not solved by any CRS. Since most CRSs incorporate comparable patch agents, we believe these failures stem not from capability gaps but from other factors: missing PoVs that never triggered patch generation, scheduling pressure from too many concurrent PoVs, and system-wide stability issues discussed earlier. One identifiable cause is deployment-time configuration trade-offs. For example, ws508 is locally patchable by MR (one agent in AT), but AT’s 30-minute per-CPV timeout—necessary to manage dozens of concurrent challenges—is largely consumed by the wireshark build alone, leaving insufficient time for the patch loop to complete.
CPVs Failed by both CRSs and MR/CC. Beyond the semantic correctness issues identified above, we examine CPVs where no valid patch was ever produced. Seven such Java cases exist: pb103, pb104, pb106, pb108, po106, po100, and po103, dominated by infinite loops (4), with the remainder being an integer overflow (pb106), a ReDoS (po106), and a JVM crash via obfuscated backdoor (po103). These bugs either lack actionable stack traces (timeouts and JVM-level crashes produce minimal localization) or involve logical complexity beyond current agent reasoning. Addressing them likely requires integrating specialized tools for specific vulnerability classes into the agent workflow, such as symbolic regex repair for ReDoS [36], rather than relying solely on generic retrieval-based approaches.
Noteworthy Challenge Projects. In shadowsocks, three CRSs patched CPVs for which they had no PoV, even though two of them require PoVs to generate patches. This occurred because all five heap-buffer-overflows (ss100–ss104) represent the same bug pattern repeated at different locations within a large JSON parsing function; some CRSs recognized the identical pattern and generated patches that fixed all instances together. In contrast, pdfbox exposes four CPVs that no team patched because CRSs never triggered them; detection-dependent pipelines cannot patch what they do not find. Finally, poi presents seven CPVs that several teams detected but none successfully patched; even CC and MR under optimized conditions produced valid fixes for only three, indicating non-trivial repair complexity.
7.5 Other Analyses
SARIF Validation. Figure 3 (v) presents SARIF validation results across 13 broadcasts (8 valid, 5 invalid). PoV-centric teams (§ 6.3) can only submit Correct when a PoV matches; thus, they were unable to assess the 5 invalid reports, leading to fewer submissions (AT: 8, FB: 7, TB: 1). With PoV evidence, FB and TB achieved 100% accuracy, but AT falsely matched PoVs to 2 invalid reports. In contrast, non-PoV-centric teams can assess all broadcasts but risk wrong answers: 42 scored 10/13 and SP 9/13.
Bundle Results.
AT TB TI FB SP 42 LC Total PoV-Patch 27/28 18/18 16/18 6/9 7/7 4/4 1/1 79/85 PoV-SARIF 1/1 – – – – – – 1/1 Patch-SARIF – – 1/1 0/2 – – – 1/3 PoV-Patch-SARIF 7/7 1/1 – 2/2 – – – 10/10 Accuracy 35/36 19/19 17/19 8/13 7/7 4/4 1/1 91/99 Score 136.2 65.1 49.8 26.4 25.3 11.0 3.2 316.9
Table VII shows bundle outcomes. Overall accuracy is high (91/99, 92%): all seven teams adopted PoV-based patch generation (§ 6.4) that naturally pairs PoVs with their patches, accounting for 86% of bundles with 93% accuracy. Patch-SARIF, the only non-PoV pairing, achieved 1/3 accuracy. Of the 8 incorrect bundles, only one is a true pairing mismatch; the other 7 failed due to unsuccessful patches, confirming patch quality as the practical bottleneck.
Resource Usage. Each team was provisioned $85K in Azure compute and $50K in LLM API credits (§ 2). Azure utilization varies widely, from 8% (LC) to 87% (AT), while LLM credit consumption is generally lower, with five of seven teams using under 25% of their LLM budget. AT, the winner, utilized the largest share of both budgets (87% of Azure, 59% of LLM credits). See Appendix K for per-provider/model breakdown.
0-Day Discovery. All seven teams discovered at least one 0-day, yielding 25 distinct vulnerabilities across 10 OSS projects, of which 12 (48%) were patched (Figure 3 (ii, iv)). See Appendix I for more details. Responsible disclosure was coordinated by Kudu Dynamics with OSTIF and ADALogics.
8 Lessons and Future Directions
The AIxCC Trifecta. AIxCC is fundamentally a test of three intertwined capabilities: research, engineering, and strategy. Around 143 hours of fully autonomous operation, CRSs tackled diverse challenges across languages, vulnerability types, and project scales—with roughly half of CPVs solvable by foundational techniques, yet many released in later phases (P3–P7) where only stable systems could capture them. Winning required balancing all three, and the finalist teams illustrate contrasting trade-offs. TI built the most agentic architecture with sophisticated technique designs, but its aggressive strategy led to the largest accuracy penalty. SP coordinated 53 components for the broadest technique coverage, but over-engineered at the expense of reliability, missing CPVs that foundational techniques could handle. AT struck the most effective overall balance, combining solid technique coverage with robust engineering and sustaining scoring across all seven phases while competitors plateaued or crashed after P3–P4. Despite these stability challenges, CRSs collectively found and fixed many CPVs that foundational techniques could not while suffering less from semantic incorrectness issues, confirming that the real bottleneck is not technique capability but robust integration into autonomous systems.
From Competition to Industry Deployment. Reflecting on the competition, we identify several areas where future efforts could further ease the path to practical deployment.
In AFC, CRSs operated on self-provisioned Azure clusters with budgets of hundreds of dollars per challenge, ensuring resource availability would not limit technical exploration. This contrasts with individual developers or small teams who need lightweight, single-machine solutions at minimal cost. Although finalist CRSs have been open-sourced, their resource usage models and runtime environments differ substantially from typical deployment settings, posing barriers to post-competition adoption. Future work could develop resource-efficient CRS variants that remain effective under constrained environments; future competitions could also introduce resource-limited tracks that account for the needs of individual developers and small teams.
Beyond resource constraints, OSS communities need time to adapt their pipelines for CRS integration, such as provisioning LLM services, standardizing CRS interfaces for broader OSS applications, defining end-to-end workflows from AI-assisted bug finding to patch submission, and ensembling multiple CRSs for combined effectiveness. Considering these needs at the design stage would lower adoption barriers for OSS maintainers and practitioners. On the post-competition side, initiatives like OSS-CRS [57] have begun to address these needs; future competitions could learn from this by co-designing deployment pathways with OSS communities for smoother transition to real-world adoption.
From Competition to Research Advancement. The competition design and team-built systems hold significant research value, yet certain design improvements could be made to further enlarge their value as research assets.
AIxCC is intrinsically a substantial experimental investment, yet its telemetry primarily serves real-time monitoring and scoring rather than retrospective analysis of why systems behaved as they did. Designing telemetry with post-hoc analysis as a first-class goal—logging intermediate outputs, decision traces, and environmental snapshots—would enable systematic studies of failure modes and technique effectiveness. Organizer-built baseline CRSs participating alongside teams would further enrich such analysis by providing reference points for comparison.
The competition’s exploration of open-source LLMs was minimal: although two teams fine-tuned open-source models, only one ultimately deployed them. The competition structure offered little incentive to invest in open-source models, as the uncertainty, cost, and data-acquisition difficulty of fine-tuning made prompt-based techniques on frontier commercial models a more predictable and cost-effective strategy. While advancing frontier AI for cybersecurity is a natural focus, open-source models offer distinct value to the OSS community through lower cost, customizability, and transparency. A dedicated sub-track comparing CRS performance under open-source models would encourage exploration in this direction.
Areas of Expansion. As the first large-scale competition of its kind, AIxCC necessarily scoped its focus. Several directions not covered in this iteration are worth exploring in future editions: full autonomy (generating harnesses and handling arbitrary build systems), multi-CRS settings (collaborative analysis or adversarial formats where CRSs attack competitors’ patches), and semantic correctness evaluation (approaches for patch semantic correctness).
9 Conclusion
AIxCC represents a milestone in autonomous cybersecurity research, demonstrating that AI-powered CRSs can discover and patch vulnerabilities in real-world software at scale. Through systematic analysis of competition design, CRS architectures, and results, we summarized technical insights of those systems, revealed both their genuine performance advances and the persistent gap between technique capability and system reliability. We hope this work serves as a foundation for future competition designs, CRS development, and practical deployment of autonomous cybersecurity systems.
10 Acknowledgment
This work is a joint effort of multiple teams and organizations involving dozens of contributors, who collectively aim to provide a systematic and insightful view of the AIxCC competition. We are grateful to every one of them for making this work possible and highlight their primary contributions.
This work was initiated and directed by Taesoo Kim, who supervised the entire research process. Cen Zhang led the structure design, the early CRS code study and team meetings, cross-team collaboration, coordination with contributors to distill the core findings of each part, and the paper drafting. Younggi Park implemented the patch analysis framework, conducted the experiments, manually validated agent-generated patches, and helped analyze and draft the patch analysis findings. Fabian Fleischer verified and integrated competition data, visualized the results (such as Figure 3), helped validate patches and drafted patch analysis, and prepared the artifact. Yu-Fu Fu compiled team statistics, studied and summarized CRS patch systems of all teams, and helped draft the patch techniques. Jiho Kim conducted PoV generation analysis in § 7.3 and contributed to the discussions in § 8. Dongkwan Kim contributed in team meeting discussions, conducted the parallel fuzzing experiments, authored § 7.5, studied teams’ bundling strategies, and helped draft § 6.4. Youngjoon Kim studied CRS SARIF techniques of all teams and some teams’ CRSs during the early code study, authored § 4, and helped draft § 6.3. Qingxiao Xu analyzed the inaccurate patch submissions across teams, and both Qingxiao Xu and Ze Sheng assisted with the analysis in § 7.5. Andrew Chin studied some CRSs’ bug finding techniques during the early code study, cross-validated the technique taxonomy in § 6, and helped check and refine the paper during the final stages. The above Team Atlanta authors, along with Hanqing Zhao, Joshua Wang, authors from Team Fuzzing Brain (Jeff Huang, Ze Sheng, and Qingxiao Xu) and Team Lacrosse (Michael Pelican, David J. Musliner), cross-validated the technique taxonomy in § 6 and conducted proofreading. Brian J. Lee helped prepare the artifact.
We sincerely thank the remaining four AIxCC teams—Trail of Bits (Michael Brown), Theori (Tyler Nighswander), Shellphish (Wil Gibbs and Yan Shoshitaishvili), and 42-b3yond-6ug (Xinyu Xing)—for their support, feedback, discussions, and meetings, as well as the internal materials they shared (e.g., Wil Gibbs shared Shellphish’s internal white paper), which helped shape our understanding of each CRS’s overall design motivations. Xinyu Xing also shared his insights and reflections on the competition, such as the gap between research and practice, during early discussions of this work.
Kudu Dynamics, as the competition organizer, provided raw materials on competition design from the organizer’s perspective, shared competition data and continually updated it as our analysis needs evolved, and held weekly meetings for communication and coordination. Nicholas Vidovich and Matthew Lehman led the organizer collaboration effort and contributed organizer-side content and review. Isaac Goldthwaite contributed content on competition rules, challenges, and scoring. Jefferson Casavant contributed content on competition infrastructure and implementation. Jon Silliman and Mikel Mcdaniel ensured the paper statistics and data aligned with the competition dataset and results. We also thank DARPA for its generous support and prompt responses to our inquiries, including granting full access to the challenge source code and competition data.
Appendix A Ethical Considerations
Stakeholder Identification. We identify three primary stakeholder groups. (1) Researchers and practitioners: the seven AFC finalist teams, competition organizers (DARPA, ARPA-H, and Kudu Dynamics), and security researchers who may build upon our findings, whose system designs, performance data, and strategic decisions are analyzed in detail. (2) Open-source community: developers and maintainers of the 24 open-source projects from which challenge projects were derived, as well as the broader OSS ecosystem that depends on the security of these projects. (3) LLM vendors: whose models were used by competing teams and whose API usage patterns are discussed.
Ethical Principles.
Beneficence. This work advances the understanding of autonomous Cyber Reasoning Systems and their application to real-world vulnerability detection. By systematically analyzing competition design, CRS architectures, and performance outcomes, we provide actionable insights that benefit the OSS security community, future competition organizers, and researchers developing autonomous security tools.
Respect for Persons. We had discussions with organizers and competing teams to gather firsthand accounts of design decisions and operational experiences, and maintained ongoing communication throughout the writing process to ensure accurate representation of their work. The AIxCC competition rules permitted publication of CRS source code, performance data, and architectural details; all finalist teams were aware that their systems and results would be subject to analysis. Teams reviewed relevant portions of the manuscript prior to submission.
Justice. All seven finalist teams are analyzed with equal rigor and presented with consistent methodology. No team is singled out or unfairly characterized. We did not re-evaluate any team’s CRS independently; all performance data originates from AFC competition results. The AFC organizers have no conflict of interest with any of the seven finalist teams and conducted fair evaluation through extensive communication and documentation. Competition data will be publicly released to ensure transparency and reproducibility, subject to DARPA’s approval and disclosure guidelines.
Respect for Law and Public Interest. The AIxCC competition operated under DARPA and ARPA-H research frameworks, with all teams agreeing to rules governing data handling and disclosure. 0-day vulnerabilities were reported through responsible disclosure processes in compliance with applicable laws. We document our methodology and data sources to enable reproducibility; competition data will be publicly released to ensure transparency and accountability.
Potential Harms. This work does not involve human subjects or private user data. Discussions with teams focused on technical methodology and system architecture, not personal or sensitive information. We identify three potential tangible harms. First, misinterpretation of performance data could cause reputational or financial harm to specific teams if rankings or analyses are taken out of context. Second, 0-day vulnerability exposure could be exploited by malicious actors before patches are available. Third, dual-use concerns arise because insights into CRS architectures could inform the development of malicious automated tools.
Mitigations. We took the following steps to address these risks. For misinterpretation of performance data, each section was cross-validated by at least two authors. We also had discussions with all seven finalist teams and maintained regular communication throughout the writing process to ensure accurate representation. For 0-day vulnerability exposure, all 0-day vulnerability data originates exclusively from the AFC competition environment. Responsible disclosure is coordinated by Kudu Dynamics with OSTIF and ADALogics under DARPA oversight, and all discovered 0-day vulnerabilities have been reported to upstream OSS maintainers. Ongoing coordination with maintainers for unpatched vulnerabilities continues through this established disclosure process. Detailed 0-day information is not included in this paper or its artifacts. For dual-use concerns, CRS performs not only vulnerability detection but also vulnerability repair. This aligns with the philosophy of OSS-Fuzz: by enabling defenders to find and patch vulnerabilities faster than malicious actors can exploit them, CRS contributes positively to the security community.
Decision to Conduct and Publish.
Decision to research. AIxCC represents the largest competition to date for LLM-based autonomous vulnerability detection and repair, yet no systematic analysis of its design, CRS approaches, or outcomes existed prior to this work. We determined that the research was justified by the need to document lessons learned, identify genuine technical advances, and surface open challenges for the security research community.
Decision to publish. The decision to publish was made after communicating with all participating teams, competition organizers, and sponsoring agencies. We believe our insights on CRS design and implementation will benefit future CRS developers and software security researchers. Since we also provide insights into vulnerability repair techniques, the benefits to the security community outweigh potential dual-use risks.
Appendix B Open Science
All data, scripts, questionnaires, and meeting notes used in this study will be released following DARPA’s official competition data release timeline.
Appendix C Scoring Details
The scoring system is centered around a developer-centric principle: reward outcomes that benefit project developers and penalize behaviors that would burden them. Organizers use this principle to determine when and to what extent should reward be given for each capability, as reflected in § 3.1:
A single PoV is not highly rewarded since it only demonstrates the vulnerability;
Patches score highest as they directly solve security issues;
SARIF assessments are useful but only offer non-core, semi-subjective report validation, therefore scoring lowest;
Bundling saves developer investigation time when correct (rewarded), but largely wastes their efforts and emotions when incorrect (penalized).
Based on those weights, organizers introduce additional scoring mechanisms, like accuracy penalty and time-decay, to ensure competition fairness and balance research incentives with practicality.
Team Score Hierarchy. A team’s total score is the sum of individual challenge scores:
Each challenge score equals the accuracy multiplier times the sum of four capability scores:
Accuracy Multiplier. The accuracy multiplier serves as a global button to encourage teams to balance effectiveness with reliability. Teams with 100% accurate submissions receive no penalty (), while lower accuracy results in a reduced multiplier. The score formula is as follows:
where the quartic formula strongly penalizes low accuracy to discourage impractical strategies, while tolerating reasonable error rates to enable research innovation. For instance, at 90% accuracy (nearly no penalty), while at 50% and at 40% (significant score reduction).
Each submission is classified as accurate, inaccurate, or neutral, where neutral submissions do not affect accuracy calculations. For a given capability, the classification rationale is to reward correct, non-duplicate submissions, penalize incorrect ones, and not penalize partially correct or duplicate submissions that provide real-world value:
PoV: One reproducible PoV per vulnerability is accurate; irreproducible PoVs are inaccurate; reproducible but duplicate PoVs are neutral, since collecting diverse PoVs can benefit patch validation in practice.
Patch: First, patches that fail to apply, build, or remediate any PoV are inaccurate. Then, the patches that fail functionality tests are neutral. This considers that CRSs may not have access to full testing functionality in practice and a CI system can help catch such issues. Finally, among passing patches, a minimal covering set is selected as accurate while the rest are inaccurate.
SARIF and Bundle: Only fully correct submissions are accurate; others are inaccurate.
Server errors and schema mismatches: neutral.
Time-Decayed Scoring Design. Teams’ submissions earn fewer points over time, with up to 50% reduction at the deadline. , , and share a common time-decay formula:
where remaining time is from submission to deadline, and total duration is the challenge window. ranges from 1.0 (immediate) to 0.5 (at deadline). The for PoV, Patch, and SARIF are 2, 6, and 1, respectively. For total duration, PoV and Patch submissions are measured from challenge start to deadline (full-mode for 12 hours while delta-mode for 6 hours), while SARIF is from broadcast time.
Note that for each vulnerability, only the last submission is scored, so earlier submissions can be superseded by later ones, allowing teams to revise their submissions with a certain penalty.
Bundle Scoring. A bundle reports pairings among PoV, Patch, and broadcast SARIF that a team identifies as addressing the same vulnerability. Unlike other capabilities, bundles can yield negative scores:
The sign is positive if all claimed pairings are correct, negative if any is incorrect. The more correct pairings, the higher the score (up to 7 points for vulnerabilities that have a SARIF broadcast, 4 for those without). CRS-generated SARIF may be included in bundles but does not affect scoring. Bundle scoring is indirectly affected by time-decay through the underlying PoV and Patch scores.
Patch Selection and Validation. Since a single patch may remediate multiple vulnerabilities, organizers must select which patches to credit. The selection algorithm identifies a minimal covering set: the smallest number of patches that collectively fix all validated vulnerabilities. When multiple patches cover the same vulnerability, specificity is preferred: patches fixing fewer vulnerabilities are chosen over broader ones, rewarding precise, targeted fixes. Patches not selected into this minimal set count as inaccurate.
To validate whether a patch truly remediates a vulnerability, organizers use all PoVs submitted by all teams, plus organizer-created ones, as test cases. A patch must remediate every PoV targeting its claimed vulnerability to be considered valid. This cross-team validation also reflects one collaboration approach among CRSs in practice.
Appendix D PoV Generation Techniques
Table VIII summarizes PoV generation techniques across teams.
AT
TB
TI
FB
SP
42
LC
Pre-Comp Corpus
Source
OSS-Fuzz; GitHub
ClusterFuzz; GitHub
ClusterFuzz; GitHub
OSS-Fuzz
Samples
Matcher
Input Format
Cov-Based
Name; Input Format
Name; Input Format
Always
LLM-Based Seed Gen
Bootstrap
Solve cov blocker
Stuck Seeds
Frontier Func
Frontier Func
LLM-Picked
Mutator/generator
Input Grammar
Testlang; libFDP [38]
Python Decoder
Nautilus [5] Grammar
Output format
Blob; Script
Script
Script
Script
Script
Blob
Semantic Feedback
LLM for IJON [6] Annot.
Improved Sanitizer
Patched (Java)
Loosened (Java)
Dict Gen
On-the-fly LLM
AFL++ [22] Dict2File; CodeQL [27]
AFL++ [22] Dict2File; Custom
Custom
Concolic Fuzzing
SymCC [48]; Custom
Directed Fuzzing
Custom distance
LLVM Slicing; WALA [65]
Parallel Fuzzing
Corpus sync
added C fuzzers
AFL++; libAFL [23]; Custom
AFL++
AFL++
AFL++
Fuzzing Pipeline
added JVM fuzzers
Custom
Bug Cand. I.D.
Cand. source †
LLM; CodeQL [27]; Sinks
LLM; Infer [39]
LLM
LLM; Entropy; CodeQL [27]; Semgrep [52]
LLM
Cand. filter
Agentic pick; Reachability
LLM confidence ranking
LLM pick
Multi-source weighted vote
Multi-LLM weighted vote
Non PoV Gen usage
PoV Gen Agent
Key context/tool †
Code; CWE; Call Path; Cov; Log; Debugger
Code; CWE
Code; Cov; Log; Debugger
Code; CWE; Call Path; Log
Code; CWE; Call Path; Cov; Log; Debugger
Main method
Iterative; ReachExploit
Iterative
Iterative; ReachExploit
Iterative
Iterative; ReachExploit
LLM-Based PoV Gen
Output format
Blob; Script
Script
Script
Script
Script
LLM PoV Gen Fuzz
Pipeline Co-op
Fuzz LLM PoV Gen
Deduplication
Stack
ClusterFuzz [29]
StackLLM classifier
Crash sigLLM
ClusterFuzz [29]
ClusterFuzz [29]
PoV hash
PoV Submission
Submission Strategy
ASAP
ASAP
ASAP
ASAP
ASAP
ASAP
ASAP
Appendix E Patch Generation Techniques
Table IX summarizes patch generation techniques across teams.
AT
TB
TI
FB
SP
42
LC
Arch Category
Multi-Arch
Multi-Agent
Multi-Agent
Single-Agent
Multi-Arch
Single-Agent
Single-Agent
Design Detail
6 standalone patching agents + Aider [26]
+ SWE-Agent [70]
RCA; Strategy; Creation; Reflection
Analyzer; Patcher; Questions
Same arch with 23 strategies:
12 Full-mode
8 Delta-mode
2 SARIF
1 Unharnessed
Triage; Programmer; Critic;
Traditional + Agentic pipelines
Test; Context; RCA; Strategy; QE
Multi-LLM workflow
Diversified Hyperparams
Temp.
Temp.
Temp.; No. of failed patches
Temp.
Agent Design
Diversified LLMs
GPT; Claude; Gemini
GPT; Claude
GPT; Claude; Gemini
Claude; GPT; Gemini
Claude; GPT
GPT; Claude; Gemini
GPT; Claude; Gemini
Standalone RCA
Multi-PoV
RCA
Up to 15 variants
Up to 3 ranked PoVs
Crash statistics (AURORA [8]) based
Non-LLM RCA
Agent with static & dynamic analysis;
ensembled ranking of multi-sources
Contextualization†
Code Indexer
ctags [14]; ast-grep [16]
tree-sitter [64]
gtags [69]
tree-sitter [64]
ctags; LSP [40]
SAST Report
Infer [39]; Joern [68]
SVF [59]; CodeQL [27]
Semgrep [52]; CodeQL [27]
CWE Guidance
40+ CWE catalog
40+ CWE repair advice
Fine-tuned
LLM
Llama [63] for contextualization
Agentic Code
Search
Dynamic Info
GDB [24]/JDB [47]
LLVM-cov [35]; JaCoCo [21]
LLVM-cov/JaCoCo
GDB/JDB
PoV Bytes
LLM Reflection
Patch Gen
No-PoV Patch
@45min; Delta diff; SAST
@50%; SAST; LLM vuln ranking
Delta diff
Validity Checks
Build
PoV Test (Gen)
Single
Up to 15/san
All
Single
Up to 20/vuln
All
Single
PoV Test
All & cross-block
Up to 15/san
All
Max 5
Up to 20/vuln
All
Single
Proj. Tests
LLM as Judge
Post-patch Fuzz
25s Fuzz (No PoV Patch Only)
5min Fuzz
Patch Valid.
Rebuild Optim.
ccache [11]; Maven [3] cache
ccache; Maven cache
Min. Patch Set Calc.
Calc. Timing
On new PoV
On new PoV
On new PoV
Every hour
Calc. Mode
Incremental; Uncovered PoVs
Recompute; All PoVs
Recompute; All PoVs
Incremental; Uncovered PoVs
Submit
Right after calc.; New patch
Right after calc.; All unsubmitted
60min; by PoV count; All unsubmitted
Right after calc.; New patch
Patch Sub.
No-PoV Patch Sub.
>45min; gated by PoV success
@50% time
@DDL-30min
Appendix F SARIF Validation Techniques
Table X summarizes SARIF validation techniques across teams.
AT TB TI FB SP 42 LC Category PoV-centric PoV-centric Bug-cand-centric PoV-centric LLM-judge-centric LLM-judge-centric LLM-judge-centric Implementation∗ LLM-Based Heuristic-Based LLM-Based Heuristic-Based; LLM-Based LLM-Based LLM-Based LLM-Based Validation Strategy Pre-validation Sanity Check† Format; File; Function; Line no. Format Format Format Format; File; Function File; Function; Line no. SARIF Report Full Report File; Function; StartLine; EndLine Function; File; Description; Rule File; StartLine; EndLine; Rule; Message Function; File; StartLine; Rule; Message Full Report Full Report PoV/Bug Cand Info Crash Log; Patch Diff (if available) StackTrace (File; Function; Line) Function; File; Description; Condition Partial Crash Log (StackTrace; Error Message) N/A Crash Log N/A Used Inputs for Validation Code Context Surrounding Context; Dynamic Retrieval‡ Enclosing Function Dynamic Retrieval Dynamic Retrieval Surrounding Context
∗ Heuristic-based follows developer-defined workflows; LLM-based relies on LLM queries.
† Verifies SARIF report format and checks whether referenced artifacts (files, functions or lines) exist before main validation.
‡ Additional code context dynamically requested by LLM.
Appendix G CP Details
LABEL:t:cp-details provides detailed information about each Challenge Project (CP) in the AIxCC final round.
Appendix H CPV Details
LABEL:t:cpv-details provides detailed information about each Challenge Project Vulnerability (CPV) in the AIxCC final round.
Appendix I 0-Day Details
LABEL:t:zeroday-details provides detailed information about 0-day vulnerabilities discovered during the competition.
Appendix J SARIF Broadcast Details
LABEL:t:sarif-broadcast-details provides detailed information about each SARIF broadcast in the AIxCC final round.
Appendix K Resource Usage
Table XI summarizes the overall resource utilization per team, Figure 4 shows the per-model token consumption, and Figure 5 shows the input-to-output token ratio per model. Token data is derived from OpenTelemetry (OTEL) logs collected by the organizers. As the collected data was incomplete due to the competition environment, the figure does not present full usage but serves as a lower-bound reference. Notably, AT exhibits the greatest model diversity, leveraging models from all four providers (OpenAI, Anthropic, Google, and xAI) as well as custom models, and most effectively utilized its allocated budgets (87% of Azure, 59% of LLM credits).
AT TB TI FB SP 42 LC Azure ($K) 73.9 18.5 20.3 63.2 54.9 38.7 7.1 OpenAI ($K) 6.6 0.6 8.2 3.1 0.4 0.4 0.4 Anthropic ($K) 20.0 20.6 3.1 7.5 2.6 0.7 0.3 Gemini ($K) 2.8 – 0.2 1.6 – – 0.1 xAI ($K) 0.1 – – – – – – All LLMs ($K) 29.4 21.1 11.5 12.2 2.9 1.1 0.6 Total ($K) 103.3 39.6 31.8 75.4 57.8 39.8 7.8
Appendix L Submission Timing
Figure 6 shows per-team submission timing. Most teams front-load within the first quarter, consistent with time-decay scoring. SARIF is the fastest (avg 5.6%) as it triggers upon broadcast; bundles are the latest (avg 35.8%) as they depend on prior submissions and are scored at the end. SP is a notable outlier, spreading submissions across the entire window.
| (a) PoV (n=2519, avg=25.0%) | (b) Patch (n=283, avg=23.8%) |
| (c) SARIF (n=48, avg=5.6%) | (d) Bundle (n=99, avg=35.8%) |
Appendix M Submission Accuracy
Table XII contains the absolute number of counted submissions (correct and incorrect, except duplicates) as well as the accuracy rate in each scoring category. Absolute submission numbers vary significantly depending on the team and category. Accuracy rates are, in general, in the range from 80%–90%, with patch rates being below average; some teams achieved 100% accuracy in certain scoring categories.
| Counted Submissions | Accuracy (%) | AM | |||||||
| Team | PoV | Patch | SARIF | Bndl | PoV | Patch | SARIF | Bndl | Penalty |
| AT | 43 | 37 | 8 | 36 | 100 | 83.8 | 75.0 | 97.2 | 0.1% |
| TB | 31 | 24 | 1 | 19 | 90.3 | 79.2 | 100 | 100 | 0.3% |
| TI | 55 | 63 | 7 | 19 | 61.8 | 31.7 | 62.5 | 89.5 | 7.2% |
| FB | 35 | 60 | 7 | 13 | 80.0 | 23.3 | 87.5 | 61.5 | 8.1% |
| SP | 31 | 11 | 12 | 7 | 90.3 | 100 | 69.2 | 100 | 0.1% |
| 42 | 45 | 4 | 13 | 4 | 91.1 | 75.0 | 76.9 | 100 | 0.3% |
| LC | 1 | 3 | 0 | 1 | 100 | 33.3 | — | 100 | 10.7% |
Appendix N CWE-Wise Performance Analysis
Figure 7 and Figure 8 show the CWE-wise performance heatmaps for PoV generation and patch generation, respectively.